# Private Fine-tuning of Large Language Models with Zeroth-order Optimization

### v2

@@ 4.3 Analysis
Different noise mechanisms for $(\varepsilon,\delta)$-DP. We now relax the privacy guarantee provided by the Laplace mechanism to approximate $(\varepsilon,\delta)$-DP. In Table 10, we compare DP-ZO instantiated with the Laplace and Gaussian mechanisms. DP-ZO (Gaussian) outperforms DP-ZO (Laplace) for strict privacy budgets such as $\varepsilon=0.5$ because it enjoys tighter accounting (Gopi et al., 2021) and lower variance (Dwork and Roth, 2014). These advantages are less significant for larger privacy budgets; for $\varepsilon=4$, the gap between DP-ZO (Gaussian) and DP-ZO (Laplace) is within $1\%$.
Our ablation study on the DP-ZO with laplace for $\varepsilon$-DP and the comparisons of Laplace and Gaussian mechanisms for $(\varepsilon,\delta)$-DP shows that DP-ZO provides a strong privacy-utility trade-off under different DP mechanisms while DP-SGD suffers from Laplace mechanisms for $(\varepsilon,\delta)$-DP, which opens the new opportunity for the synergy between DP mechanisms and large language models.
@@ 5 DISCUSSION
In Section 4.2 we showed that DP-ZO obtains competitive privacy-utility tradeoff. Now we examine the amount of engineering effort necessary to scale DP-SGD to larger models, a topic on which many papers have been written (Bu et al., 2023b, c; Yousefpour et al., 2021; Li et al., 2022b; He et al., 2023; Bu et al., 2023a). We find that DP-ZO seamlessly scales to larger models and believe its simplicity presents a compelling alternative to DP-SGD for practitioners.
@@
DP-SGD. Differentially Private Stochastic Gradient Descent (DP-SGD) (Song et al., 2013; Abadi et al., 2016) is the standard privacy-preserving algorithm to train models on private data, with an update rule given by $w^{(t+1)}=w^{(t)}-\frac{\eta_{t}}{|B|}\left(\sum_{i\in B}\frac{1}{\text{c}}\texttt{clip}_{\text{c}}(\nabla\ell(x_{i},w^{(t)}))+\sigma\xi\right)$ where the changes to SGD are the per-sample gradient clipping $\texttt{clip}_{\text{c}}(\nabla\ell(x_{i},w^{(t)}))=\frac{c\times\nabla\ell(x_{i},w^{(t)})}{\max(c,||\nabla\ell(x_{i},w^{(t)})||_{2})}$ and addition of noise sampled from a $d$-dimensional Gaussian distribution $\xi\sim\mathcal{N}(0,1)$ with standard deviation $\sigma$. DP-SGD is the marquee algorithm for privacy-preserving machine learning, but it requires implementing per-example gradient clipping. This creates a slew of challenges for deploying DP-SGD.
@@
Computational and memory challenges in DP-SGD. DP-SGD requires the computation of per-example gradients, which can be naively implemented by storing each gradient in the batch separately. This approach inflates the memory overhead by a factor of $B$, where $B$ is the batch size. Tensorflow Privacy avoids this issue by clipping microbatches rather minibatches, which does not slow down training but increases the noise added and therefore hurts utility. Jax can automatically vectorize the per-sample gradient computation, but training is still slowed down. Recently, specialized libraries have been developed that instead analytically compute the norm of the gradients for different layers (Li et al., 2022b; Bu et al., 2023d; Ding et al., 2024). This requires actually implementing the computation, which is challenging for new layers. Parameter efficient fine-tuning methods (Yu et al., 2022; Bu et al., 2024) can help reduce such computation cost by reducing the number of trainable parameters. However, as discussed in Section 4.3, those methods still incur much more memory cost than DP-ZO for long sequences. Besides, when model cannot be loaded into a single GPU, model parallelism is needed to load large models across several GPUs and per-example gradient norm clipping requires additional implementation, both for gradient clipping and communication across device. He et al. (2023) investigate how to make group-wise gradient clipping efficient and achieve good performance in DP-SGD and fine-tunes 175B GPT3 model with 16 V100 GPUs each with 32 gigabytes of VRAM. Bu et al. (2023a) implements DP-SGD based on Zero Redundancy Optimizer (Rajbhandari et al., 2020) to scale model size up to GPT-100B and maintain efficiency. It currently also only supports layer-wise or block-wise clipping.
5.1 DP-ZO Scales Seamlessly
@@
Model parallelism in DP-ZO. To train large models like OPT-66B, whose parameters cannot be loaded into memory on a single A100 GPU, we need to implement some form of parallelism across GPUs. It is easy for such parallelism in DP-ZO (in the simplest form, just running DP-ZO on a machine with 2 GPUs will prompt HuggingFace to implement naive model parallelism), while much more effort in DP-SGD as we previously discussed.
@@
Data parallelism in DP-ZO. To synchronize model state between GPUs in data-parallel-DP-ZO, we just transfer the random seed and its corresponding half-precision float16 scalar step size; this is just a few bytes. However, first-order approaches such as DP-SGD require the transfer of gradients across devices to update all the models, necessitating expensive allgather and reduce operations. This communication overhead is $1.5d$ in PyTorch FSDP, where $d$ is the size of the model.
DP-ZO does not store gradients. DP-ZO does not store activations or gradients in the forward pass, thus conserving memory. DP-SGD needs to store the activations at each GPU under pipeline parallelism to clip the per-example gradient (He et al., 2023), which will fill up the GPU memory and limit new microbatches from being processed.
@@
DP-ZO is storage and communication-efficient even after training has completed. DP-ZO offers significant advantages in terms of storage and communication efficiency, especially beneficial for bandwidth-constrained environments like edge devices. Unlike traditional methods where the difference in model parameters $\theta_{0}-\theta_{f}$ is shared—which could amount to multiple gigabytes for large models—DP-ZO allows for the storage and transmission of a sequence of updates. This sequence is represented as an array of tuples $[(\textsc{seed}_{0},0.54),\cdots,(\textsc{seed}_{f},-0.14)]$, where each tuple contains a seed and a step size, taking up only 4 bytes. Even for $1\times 10^{4}$ fine-tuning iterations, this array would require less than 1MB of storage, representing a substantial reduction in both storage and communication overhead. We can apply these weight differences to a model by simply iterating over the array, sampling from the PRNG using the given seed, scaling that random vector, and applying it to the current model parameters. This procedure is highly efficient, as it involves only sequential memory accesses and scalar floating-point operations.
6 RELATED WORK
Privacy preserving large language models. Recent studies have leveraged DP-SGD to fine-tune large language models. Li et al. (2022b) provide methods for fine-tuning large language models with DP-SGD by ghost clipping to mitigate the memory burden of per-sample gradient clipping. Yu et al. (2022) report compelling results by only updating a sparse subset of the LLMs with parameter efficient fine-tuning (PEFT) methods such as LoRA (Hu et al., 2022). He et al. (2023) leverage group-wise clipping with adaptive clipping threshold and privately fine-tune the 175 billion-parameter GPT-3. Duan et al. (2023); Li et al. (2022b) also consider private prompt tuning by adding noise to the soft prompt (Li and Liang, 2021; Lester et al., 2021). Du et al. (2023) add non-i.i.d. noise from a matrix Gaussian distribution to directly perturb embedding in the forward pass of language models. With the emergence in-context learning of large language models (Brown et al., 2020), recent works (Duan et al., 2023; Wu et al., 2024; Tang et al., 2023b) study private in-context learning of large language models without fine-tuning.
Private zeroth-order optimization. Most recently, a concurrent work (Zhang et al., 2024a) also considers the same DP-SPSA algorithm for zeroth-order optimization. Our method and Zhang et al. (2024a) are functionally the same up to constants, and our work focuses on an empirical evaluation of the method, whereas Zhang et al. (2024a) extends the convergence analysis of Malladi et al. (2023) to DP as shown in Appendix B of Zhang et al. (2024a). There is a slight difference for the generation of random perturbation of Zhang et al. (2024a) and our Algorithm 1. Zhang et al. (2024a) uses the random unit vector for the perturbation and the convergence analysis is based on such set-up, whereas our perturbation is a normally distributed vector. Note that Algorithm 1 and 2 in Zhang et al. (2024a) also scales the unit vector by the square root of the model dimension, so the two approaches are functionally the same. We reimplemented our perturbation method based on the algorithms in Zhang et al. (2024a), and we obtain $82.32_{0.82}$ for fine-tuning OPT-13B by LoRA finetuning on SQuAD with $(1,10^{-5})$-DP, that is comparable as our main result in Table 1. Besides, Our work provides a systematic study of DP-ZO and DP-SGD on privacy-utility trade-off including different models and different noise mechanisms, as well as computation cost and privacy implication.
@@ Appendix A Privacy Analysis
Proposition A.1 (Basic Composition theorem [Dwork and Roth, 2014]).
@@
Proposition A.2 (Privacy Amplification via Subsampling [Balle et al., 2018]).
@@
Gaussian Mechanism
As outlined in Line 11, the $\ell_{2}$ sensitivity of  Algorithm 1 is $\mathrm{C}$ and we are adding $\mathcal{N}(0,\mathrm{C}^{2}\sigma^{2})$ noise to the estimated loss. We analyze the composition of subsampled Gaussians with the privacy loss variable accounting approach of Gopi et al. [2021].
Laplace Mechanism
@@
Laplace mechanism can gives a pure DP guarantee of $\delta=0$ which can be of interest in some scenarios. Here we first analyze the pure $\varepsilon$-DP guarantee provided by laplace mechanism and then provide the analysis for approximate $(\varepsilon,\delta)$-DP analysis.
@@
Pure $\varepsilon$-DP by Laplace mechanism. We use data in a single batch instead of all training data to compute the gradients in each updates. For the privacy analysis for Laplace mechanism in Algorithm 1, when we sample each batch in the poisson manner, we could leverage Proposition A.2 to compute the private amplification by subsampling. We first analyze the privacy cost for one step by the Laplace mechanism. At each step, we sample a new batch of data with the sample rate of $p=B/|\mathcal{D}|$. As outlined in 11, the $\ell_{1}$ sensitivity of Algorithm 1 is $\mathrm{C}$. By Proposition 2.3, the privacy cost at one step would cost $(1/\sigma,0)$-DP on this batch. By Proposition A.2, the privacy cost at one step would cost $(\log(1+p\cdot(e^{1/\sigma}-1)),0)$-DP on the full dataset $\mathcal{D}$. By Proposition A.1, the privacy cost of Algorithm 1 instantiated with Laplace mechanism satisfies $(T\cdot\log(1+p\cdot(e^{1/\sigma}-1)),0)$-DP.
@@
Approximate $\varepsilon$-DP by Laplace mechanism. We can also get tighter composition of $\varepsilon$ with relaxation to $\delta>0$. The most straight forward way is to instantiate the PRV of random response with $(\log(1+p\cdot(e^{1/\sigma}-1)),0)$ because the dominating pair for random response is a dominating pair for the pure DP mechanism. Then, we can use the numerical composition of Pure DP PRV by Gopi et al. [2021]. Note that this method is agnostic to the DP mechanisms used for pure $\varepsilon$-DP. We now provide a more fine-grained privacy analysis for the laplace mechanism. Specifically, we could compute the privacy cost of composition for the Laplace mechanism by Monte Carlo based DP accountant [Wang et al., 2023]. Note that since we are dealing with scalar values, our mechanism in each iteration will be a one dimensional Laplace mechanism. Let $b$ be the scale of Laplace noise, $p$ the sub-sampling rate, and assume the sensitivity is $1$, and assume we are doing composition for $T$ iterations, each iteration with sampling rate $p$. By Zhu et al. [2022] we know that the pair of distribution $(P,Q)$ dominating pair for a single dimensional Laplace mechanism, where $P$ and $Q$ are distributed according to the following pdfs,
Therefore, ($P,(1-p)\cdot P+p\cdot Q$) is the dominating pair for the sub-sampled Laplace. We plug this into the standard Monte-Carlo accountant of Wang et al. [2023] (without importance sampling, see Algorithm 2 and Theorem 10 in Wang et al. [2023]) while using $10^{10}$ samples to calculate the $\delta$ at a given value of $\epsilon$. Also, using the analytical accountant explained above, we always make sure that $\mathbb{E}[\hat{\delta}_{MC}^{2}]$ is bounded by $10^{-8}$ (We use the fact that $\mathbb{E}[\hat{\delta}_{MC}^{2}]$ is bounded by $\mathbb{E}[PRV^{2}]$ and the fact the $PRV$ is always bounded for Laplace mechanism.). This ensures that the error of our estimation of $\delta$ is at most $10^{-8}$ with probability at least $1-10^{-5}$. Putting all together, for all reported values of $\epsilon$, our $\delta$ is bounded by $10^{-5}$, with probability at least $0.99999$. This privacy analysis is tighter with $\varepsilon$ is high compared to the former analysis which uses the pure DP PRV accountant. This is consistent with the intuition. As we increase the distance between the Laplace dominating pairs, the probability of sampling points from the area between the centers increases. And that is where the Laplace Mechanism is different from the Randomized Response. We present the accounting results for the Laplace method to achieve $(\varepsilon,\delta)$-DP by these two accounting methods in Table 12. Table 12 shows that the Monte Carlo based DP accountant can give tighter analysis for the Laplace mechanism for $(\varepsilon,\delta)$-DP than the pure $\varepsilon$-DP PRV method.
Table 11: $\varepsilon$-DP by Laplace. BSZ=20, Steps=2000.
@@
Table 12: $(\varepsilon,\delta)$-DP guarantee for Laplace. $\delta=10^{-5}$. $|\mathcal{D}|=1000$. BSZ=16, Steps=75000.

### v3

@@ 4.3 Analysis
Table 11: (DP-)ZO and (DP-)SGD across different number of training samples $n$. $(3,4.7\times 10^{-6})$ for QNLI on RoBERTa-base. $(1,10^{-5})$ for SST2 on OPT-13B. For full set, $n=104743$ for QNLI and $n=66849$ for SST2.
QNLI SST2
DP-ZO ZO DP-SGD SGD DP-ZO ZO DP-SGD SGD
$n=1000$ $76.19$ $79.85$ $73.27$ $84.40$ $91.19$ $93.69$ $90.25$ $95.64$
full set $79.85$ $79.99$ $87.40$ $89.49$ $93.00$ $94.84$ $95.07$ $96.33$
Besides the few-shot setting, we now scale the size of training data size up to more than 100k samples and compare the performance of DP-SGD and DP-ZO. For the ease of comparison, we follow Li et al. (2022b); Yu et al. (2022) and conduct experiments for DP-ZO and DP-SGD on QNLI with RoBERTa-base with a range of examples at $(3,4.7\times 10^{-6})$-DP and report result in Table 10. Similar to previous observation, Table 10 shows that at a small data regime, i.e., number of samples $n=1000$, DP-ZO achieves comparable performance as DP-SGD. DP-ZO can also improve its performance within more samples. Compared to DP-SGD, we notice that when increasing the data size $n$, there is a utility gap between DP-ZO and DP-SGD. In addition, we compare the utility of DP-ZO, ZO, DP-SGD, SGD under different $n$ in Table 11 to further understand the utility gap between DP-ZO and DP-SGD. Due to the limited computation resources, we provide results for RoBERTa-base on QNLI and OPT-13B (LoRA fine-tuning) on SST2 in Table 11 for $n=1000$ and full training set. We observe a similar utility gap between DP-ZO and DP-SGD as QNLI in Table 10 for SST2 in Table 11 when increasing $n$ to the full training set. The utility gap of SGD and ZO on QNLI and SST2 in Table 11 indicate that the challenge of utility for ZO when the number of training samples $n$ increases. Therefore there is also the utility gap between DP-ZO and DP-SGD when $n$ increases. This limitation shows an open problem on how to improve data efficiency in zeroth-order optimization (Zhang et al., 2024c; Zhao et al., 2024) and DP-ZO, and we leave this data efficiency improvement in ZO as future work.
Different noise mechanisms for $(\varepsilon,\delta)$-DP. We now relax the privacy guarantee provided by the Laplace mechanism to approximate $(\varepsilon,\delta)$-DP. In Table 10, we compare DP-ZO instantiated with the Laplace and Gaussian mechanisms. DP-ZO (Gaussian) outperforms DP-ZO (Laplace) for strict privacy budgets such as $\varepsilon=0.5$ because it enjoys tighter accounting (Gopi et al., 2021) and lower variance (Dwork & Roth, 2014). These advantages are less significant for larger privacy budgets; for $\varepsilon=4$, the gap between DP-ZO (Gaussian) and DP-ZO (Laplace) is within $1\%$.
Our experiments on the DP-ZO with laplace for $\varepsilon$-DP and the comparisons of Laplace and Gaussian mechanisms for $(\varepsilon,\delta)$-DP shows that DP-ZO provides a strong privacy-utility trade-off under different DP mechanisms while DP-SGD suffers from Laplace mechanisms for $(\varepsilon,\delta)$-DP, which opens the new opportunity for the synergy between DP mechanisms and large language models.
@@ 5 DISCUSSION
Discussion of our work and DPZero (Zhang et al., 2024a). Most recently, a concurrent work (Zhang et al., 2024a) also privatizes the SPSA algorithm in zeroth-order optimization method. Zhang et al. (2024a) propose DPZero and provide the convergence guarantee for DPZero that is independent of the model dimension in private training. Zhang et al. (2024a) presented experimental results for DPZero/DP-SGD on Roberta-large model and DPZero on OPT 1.3B-6.7B models. Our work focus on understanding the privacy-utility trade-off for DP-ZO as well as the privacy implication and the practical implication of DP-ZO. We provide a comprehensive understanding of the privacy-utility trade-off for DP-ZO and DP-SGD for OPT-1.3B to 66B models. We consider DP-ZO as a general framework for private machine learning. We found that DP-ZO is compatible with Laplace mechanism and enabling $\varepsilon$-DP with a reasonable privacy-utility trade-off while DP-SGD with Laplace mechanism suffers to optimize. The result indicates that DP-ZO can open the new opportunity for the synergy between DP mechanisms and large language models. To the best of our knowledge, we provide the first experimental result that measures the privacy leakage in (DP-)ZO. We also conduct experiments for analyzing the effect of data size and the memory consumption for limited resource scenarios. Our findings could provide insights on future directions for improving the analysis and implementation of DP-ZO methods.
@@
Discussion on DP-SGD and DP-ZO. In Section 4.2 we showed that DP-ZO obtains competitive privacy-utility tradeoff. Now we examine the amount of engineering effort necessary to scale DP-SGD to larger models, a topic on which many papers have been written (Bu et al., 2023b; c; Yousefpour et al., 2021; Li et al., 2022b; He et al., 2023; Bu et al., 2023a). We find that DP-ZO seamlessly scales to larger models and believe its simplicity presents a compelling alternative to DP-SGD for practitioners.
@@
DP-SGD. Differentially Private Stochastic Gradient Descent (DP-SGD) (Song et al., 2013; Abadi et al., 2016) is the standard privacy-preserving algorithm to train models on private data, with an update rule given by $w^{(t+1)}=w^{(t)}-\frac{\eta_{t}}{|B|}\left(\sum_{i\in B}\frac{1}{\text{c}}\texttt{clip}_{\text{c}}(\nabla\ell(x_{i},w^{(t)}))+\sigma\xi\right)$ where the two changes to SGD are the per-sample gradient clipping $\texttt{clip}_{\text{c}}(\nabla\ell(x_{i},w^{(t)}))=\frac{c\times\nabla\ell(x_{i},w^{(t)})}{\max(c,||\nabla\ell(x_{i},w^{(t)})||_{2})}$ and addition of noise sampled from a $d$-dimensional Gaussian distribution $\xi\sim\mathcal{N}(0,1)$ with standard deviation $\sigma$. DP-SGD is the marquee algorithm for privacy-preserving machine learning, but it requires implementing per-example gradient clipping. This creates a slew of challenges for deploying DP-SGD.
@@
Computational and memory challenges in DP-SGD. DP-SGD requires the computation of per-example gradients, which can be naively implemented by storing each gradient in the batch separately. This approach inflates the memory overhead by a factor of $B$, where $B$ is the batch size. TensorFlow Privacy avoids this issue by clipping microbatches rather minibatches, which does not slow down training but increases the noise added and therefore hurts utility. Jax can automatically vectorize the per-sample gradient computation, but training is still slowed down. Recently, specialized libraries have been developed that instead analytically compute the norm of the gradients for different layers (Li et al., 2022b; Bu et al., 2023d; Ding et al., 2024). This requires actually implementing the computation, which is challenging for new layers. Parameter efficient fine-tuning methods (Yu et al., 2022; Bu et al., 2024) can help reduce the computation cost by reducing the number of trainable parameters. However, as discussed in Section 4.3, those methods still incur much more memory cost than DP-ZO for long sequences. Besides, when model cannot be loaded into a single GPU, model parallelism is needed to load large models across several GPUs and per-example gradient norm clipping requires additional implementation, both for gradient clipping and communication across device. He et al. (2023) investigate how to make group-wise gradient clipping efficient and achieve good performance in DP-SGD and fine-tunes 175B GPT3 model with 16 V100 GPUs each with 32 gigabytes of VRAM. Bu et al. (2023a) implements DP-SGD based on Zero Redundancy Optimizer (Rajbhandari et al., 2020) to scale model size up to GPT-100B and maintain efficiency. It currently also only supports layer-wise or block-wise clipping.
@@
Model parallelism and data parallelism in DP-ZO. To train large models like OPT-66B, whose parameters cannot be loaded into memory on a single A100 GPU, we need to implement some form of parallelism across GPUs. It is easy for such parallelism in DP-ZO (in the simplest form, just running DP-ZO on a machine with 2 GPUs will prompt HuggingFace to implement naive model parallelism), while much more effort in DP-SGD. To synchronize model state between GPUs in data-parallel-DP-ZO, we just transfer the random seed and its corresponding half-precision float16 scalar step size; this is just a few bytes. However, first-order approaches such as DP-SGD require the transfer of gradients across devices to update all the models, necessitating expensive allgather and reduce operations. This communication overhead is $1.5d$ in PyTorch FSDP, where $d$ is the size of the model.
@@
DP-ZO is storage and communication-efficient even after training has completed. DP-ZO offers significant advantages in terms of storage and communication efficiency, especially beneficial for bandwidth-constrained environments like edge devices (Li et al., 2024). Unlike traditional methods where the difference in model parameters $\theta_{0}-\theta_{f}$ is shared which could amount to multiple gigabytes for large models, DP-ZO allows for the storage and transmission of a sequence of updates. This sequence is represented as an array of tuples $[(\textsc{seed}_{0},0.54),\cdots,(\textsc{seed}_{f},-0.14)]$, where each tuple contains a seed and a step size, taking up only 4 bytes. Even for $1\times 10^{4}$ fine-tuning iterations, this array would require less than 1MB of storage, representing a substantial reduction in both storage and communication overhead. We can apply these weight differences to a model by simply iterating over the array, sampling from the PRNG using the given seed, scaling that random vector, and applying it to the current model parameters. This procedure is highly efficient, as it involves only sequential memory accesses and scalar floating-point operations.
DP-SGD vs. DP-ZO. We have discussed the memory efficiency of DP-ZO over DP-SGD in details. This is helpful in limited memory scenarios as we show in Figure 5. While being much more memory efficient, DP-ZO achieves comparable performance as DP-ZO on several tasks including OPT-13B models on $1000$ SQuAD examples with Gaussian mechanism (Table 1) and provides reasonable privacy-utility trade-off for $\varepsilon$-DP (Table 3), which is promising. We also note that the utility challenge of DP-ZO when scaling with more samples (Table 8), mostly due to the gap of ZO and SGD (Table 11). This indicates the future improvement for data efficiency in zeroth-order optimization, that will benefit DP-ZO with Gaussian and Laplace mechanism.
@@ Appendix A Privacy Analysis
We first introduce two propositions for DP analysis.
Proposition A.1 (Basic Composition theorem (Dwork & Roth, 2014)).
@@
Proposition A.2 (Privacy Amplification via Subsampling (Balle et al., 2018)).
@@
Analyzing the privacy guarantees of a specific mechanism can be done via worst-case inputs to a mechanism $M$ leading to a pair of worst-case distributions (Dwork & Roth, 2014). Meiser & Mohammadi (2018) introduce a novel method to approximately compose the privacy curves based on the discretized version of privacy loss random variables (Dwork & Rothblum, 2016), whose distribution is called privacy loss distribution (PLD). Meiser & Mohammadi (2018) provide the worst-case pair of distributions for basic mechanisms such as Gaussian mechanism and Laplace mechanism. Sommer et al. (2019) derive the exact analytical and closed formula for the Gaussian mechanism. Zhu et al. (2022) formalize a rigorous notion of the “worst-case” PLD for DP mechanism under the name dominating PLDs. Therefore, following Zhu et al. (2022), we introduce the following definitions.
Definition A.3 (Hockey-stick Divergence).
For $\alpha>0$, the Hockey-stick divergence is defined as $H_{\alpha}(P\|Q):=\mathbb{E}_{o\sim Q}[(\frac{P(o)}{Q(o)}-\alpha)_{+}],$ where $(x)_{+}:=x\mathbf{1}(x\geq 0)$.
Lemma A.4.
(Barthe et al., 2016) For a randomized algorithm $\mathcal{M}$, $\sup_{D\simeq D^{\prime}}H_{e^{\varepsilon}}(\mathcal{M}(D)\|\mathcal{M}(D^{\prime}))\leq\delta$ is equivalent to Definition 1.
We follow Zhu et al. (2022) and formalize $\delta$ as a function of $\varepsilon$.
Definition A.5 (Optimal Privacy Curve).
The optimal privacy curve of a mechanism $\mathcal{M}$ is the function $\delta_{\mathcal{M}}:\mathbb{R}^{+}\rightarrow[0,1]$ s.t. $\delta_{\mathcal{M}}(\varepsilon):=\sup_{D\simeq D^{\prime}}H_{e^{\varepsilon}}(\mathcal{M}(D)\|\mathcal{M}(D^{\prime}))$.
Definition A.6 (Dominating Pair of Distributions (Zhu et al., 2022)).
$(P,Q)$ is a dominating pair of distributions for $\mathcal{M}$ (under neighboring relation $\simeq$) if for all $\alpha\geq 0$
$\sup_{D\simeq D^{\prime}}H_{\alpha}(\mathcal{M}(D)\|\mathcal{M}(D^{\prime}))\leq H_{\alpha}(P\|Q)$ (3)
When $P,Q$ is chosen such that equation 3 takes “$=$” for all $\alpha$, $(P,Q)$ is a tight dominating pair of distributions or simply, tightly dominating.
Based on Sommer et al. (2019); Koskela et al. (2020); Gopi et al. (2021), we follow Zhu et al. (2022) and define the privacy loss random variable as follows.
Definition A.7 (Privacy Loss Random Variable).
We call $Y:=\log\frac{P(o)}{Q(o)},o\sim P$ the privacy loss random variable (PRV) for mechanism $\mathcal{M}$ associated with dominating pair $(P,Q)$.
Zhu et al. (2022) prove that any privacy mechanism has a tightly dominating pair of distributions and provide the dominating pairs for basic privacy mechanisms such as Gaussian mechanism and Laplace mechanism.
Further, based on Definitions A.3 and A.7, $H_{e^{\varepsilon}}(P\|Q)$ can be written as an expectation: $H_{e^{\varepsilon}}(P\|Q)=\mathbb{E}_{Y}\left[\left(1-e^{\varepsilon-Y}\right)_{+}\right]$. $\delta_{\mathcal{M}}(\varepsilon)$ can be bounded by first identifying $\mathcal{M}$’s dominating pair distributions as well as the associated PRV $Y$, and then computing this expectation, denoted as $\delta_{Y}(\varepsilon):=\mathbb{E}_{Y}\left[\left(1-e^{\varepsilon-Y}\right)_{+}\right]$. $\delta_{\mathcal{M}}(\varepsilon)\leq\delta_{Y}(\varepsilon)$, where $\delta_{\mathcal{M}}(\varepsilon)=\delta_{Y}(\varepsilon)$ for tight dominating pair $(P,Q)$.
As we sample a batch of samples from the training data via poisson subsampling in each iteration for experiments, we introduce the subsampling amplification for dominating pair following Zhu et al. (2022).
Proposition A.8.
(Zhu et al., 2022) Let $\mathcal{M}$ be a randomized algorithm. If $(P,Q)$ dominates $\mathcal{M}$ then $(P,(1-p)P+pQ)$ dominates $\mathcal{M}\circ S_{\textbf{{Poisson}}}^{p}$ for add neighbors and $((1-p)Q+pP,Q)$ dominates $\mathcal{M}\circ S_{\textbf{{Poisson}}}^{p}$ for remove neighbors , where $S_{\textbf{Poisson}}^{p}$ is poisson sampling with subsampling rate $p$.
Given a dominating pair, Zhu et al. (2022) show that if the characteristic function of the PLDs has an analytical expression, for example, Gaussian mechanism, the tight DP composition bounds can be computed in $O(1)$ times. However, without the closed form, such as the sub-sampled Gaussian and sub-sampled Laplace, numerical accounting methods for $T$ steps are proposed to approximate the integral composition formula and compute upper bounds on the privacy budget parameters $(\varepsilon,\delta)$ in DP (Koskela et al., 2020; Koskela & Honkela, 2021; Koskela et al., 2021; Gopi et al., 2021; Doroshenko et al., 2022; Alghamdi et al., 2023; Wang et al., 2023). There are multiple open-source implementations for numerically accurate composition including Microsoft (2021) and Google’s-DP-Library (2020) that are used by DP-SGD experiments (Yu et al., 2022; Chua et al., 2024) .
We now provide the privacy analysis for DP-ZO.
Gaussian Mechanism.
As outlined in Line 10, the $\ell_{2}$ sensitivity of  Algorithm 1 is ${C}$ and we are adding $\mathcal{N}(0,{C}^{2}\sigma^{2})$ noise to the estimated loss. From Zhu et al. (2022), we know that the tight dominating pair for Gaussian mechanisms is $P=\mathcal{N}(0,\sigma^{2})$ and $Q=\mathcal{N}(1,\sigma^{2})$. By Proposition A.8, each step therefore satisfies $(\varepsilon,\delta)$-DP where
$\text{max}(H_{e^{\varepsilon}}(\mathcal{N}(0,\sigma^{2})\|(1-p)\mathcal{N}(0,\sigma^{2})+p\mathcal{N}(1,\sigma^{2})),H_{e^{\varepsilon}}((1-p)\mathcal{N}(1,\sigma^{2})+p\mathcal{N}(0,\sigma^{2})\|\mathcal{N}(1,\sigma^{2})))\leq\delta$ (4)
For $T$ step, we analyze the compositions of subsampled Gaussian for $T$ steps with the PRV accountant of Gopi et al. (2021). We provide the privacy parameters we use for Gaussian mechanism in Table 16 and Table 17. We also provide the example code for using Microsoft PRV Accountant library (Microsoft, 2021) below.
1from prv_accountant import PoissonSubsampledGaussianMechanism
2from prv_accountant import PRVAccountant
3sample_rate = 16/1000
4sigma = 16.4
5step = 75000
6delta = 1/(1e5)
7mech = PoissonSubsampledGaussianMechanism(
8 noise_multiplier=sigma,
9 sampling_probability=sample_rate,)
10accountant = PRVAccountant(
11 prvs=[mech],
12 max_self_compositions=[step+10],
13 eps_error=0.001,
14 delta_error=1e-10)
15eps_low, eps_est, eps_up = accountant.compute_epsilon(delta=delta, num_self_compositions=[step])
16print(eps_low, eps_est, eps_up)
Listing 1: DP accounting using Microsoft PRV Accountant (Microsoft, 2021). We ensure the upper bound $eps_{up}$ does not exceed the desired $\varepsilon$ at the specified $\delta$.
Laplace Mechanism.
@@
Laplace mechanism can give a pure DP guarantee of $\delta=0$ which can be of interest in some scenarios. Here we first analyze the pure $\varepsilon$-DP guarantee provided by Laplace mechanism and then provide the analysis for approximate $(\varepsilon,\delta)$-DP analysis.
@@
Pure $\varepsilon$-DP by Laplace mechanism. We use data in a single batch instead of all training data to compute the gradients in each updates. For the privacy analysis for Laplace mechanism in Algorithm 1, when we sample each batch in the poisson manner, we could leverage Proposition A.2 to compute the private amplification by subsampling. We first analyze the privacy cost for one step by the Laplace mechanism. At each step, we sample a new batch of data with the sample rate of $p=B/|\mathcal{D}|$. As outlined in 10, the $\ell_{1}$ sensitivity of Algorithm 1 is ${C}$. By Proposition 2.3, the privacy cost at one step would cost $(1/\sigma,0)$-DP on this batch. By Proposition A.2, the privacy cost at one step would cost $(\log(1+p\cdot(e^{1/\sigma}-1)),0)$-DP on the full dataset $\mathcal{D}$. By Proposition A.1, the privacy cost of Algorithm 1 instantiated with Laplace mechanism satisfies $(T\cdot\log(1+p\cdot(e^{1/\sigma}-1)),0)$-DP.
@@
Table 12: Privacy parameters for Table 3 and Table 8. $\varepsilon$-DP by Laplace. BSZ=20, Steps=2000.
@@
Table 13: Privacy parameters for Table 10. $(\varepsilon,\delta)$-DP guarantee for Laplace. $\delta=10^{-5}$. $|\mathcal{D}|=1000$. BSZ=16, Steps=75000.